OpenStudy (anonymous):

COLLEGE PHYSICS QUESTION: Every few years, winds in Boulder, Colorado, attain sustained speeds of 45.0 m/s (about 100 mi/h) when the jet stream descends during early spring. Approximately what is the force due to the Bernoulli effect on a roof having an area of 245 m2? Typical air density in Boulder is 1.14 kg/m3 , and the corresponding atmospheric pressure is 8.89 × 104 N/m2 . (Bernoulli’s principle assumes a laminar flow. Using the principle here produces only an approximate result, because there is significant turbulence.) please help!

OpenStudy (surry99):

ok lets start with the Bernoulli equation...what does is it?

OpenStudy (anonymous):

\[P _{1}+1/2\rho v ^{2}=P _{2}+1/2\rho v ^{2}\]

OpenStudy (surry99):

how would you calculate the static pressure associated with the air moving at 45 m/s?

OpenStudy (anonymous):

... I have no idea. Im so lost. I know that pressure is Force/area. but what is static pressure?

OpenStudy (surry99):

ok..if we assume the structure is closed (like a house) the air inside the house is stationary and has a pressure = to the atmospheric pressure. However, the moving air outside has a loer pressure. Lets say the air inside is 1 and the air outside is 2. So For 1, P1 = 8.89 *10^4 N/m^2 v1 = 0 (the air is essentially not moving For 2 P2 = ? V2 = 45 m/s Now the density is assume to be constant for both 1 and 2. Can you now solve for P2?

OpenStudy (surry99):

In the Bernoulli equation we refer to P1 and P2 as static pressures. The 1/2*rho*v^2 terms are called the dynamic pressures.

OpenStudy (anonymous):

is the pressure in the house like regular atm pressure, or the atmospheric pressure that's provided?

OpenStudy (surry99):

It is the value provided in the problem which accounts for the fact Boulder is at altitude

OpenStudy (anonymous):

ok. so then if the densities are constant can they be canceled out?

OpenStudy (surry99):

no, plug them in and you will see they can not be cancelled

OpenStudy (anonymous):

I think im getting it. so is meters/second the correct unit to use In the equation? for v?

OpenStudy (surry99):

yes

OpenStudy (anonymous):

ok so I set it up as 8.89*10^4+.5(1.14)(0^2m/s)=P2+.5(1.14)(45^2m/s) right?

OpenStudy (surry99):

yes

OpenStudy (surry99):

remember the m/s units are also squared...(45 m/s) ^2

OpenStudy (anonymous):

and then I multiply by the area to find the force?

OpenStudy (surry99):

not quite...P2 is the pressure pushing down on the roof but P1 is pushing up. Since P1 is greater than P2 , the net force is up so: (P1 - P1) * Area = net force.... draw a FBD to convince yourself this is coorect.

OpenStudy (anonymous):

no I understand, it just didn't ask for net force, but I get it.

OpenStudy (anonymous):

**(P1-P2)*Area =)

OpenStudy (surry99):

Awesome!

OpenStudy (surry99):

oops for the typo...yes (P1 - P2) *area

OpenStudy (anonymous):

yayyyyy I got it right ^-^ F=282791.25 N thank you!!

OpenStudy (surry99):

you are very welcome!